# HidePrint: Hiding the Radio Fingerprint via Random Noise

### v1

@@ 8 Comparison with other RFF solutions
The RFF methodology used in the previous section to validate HidePrint involves the pre-processing of IQ samples into images, as described in Sect. 6. This strategy has been proven to be effective in mitigating the impact of noise on the RFF process [8]. To provide further insights and show that HidePrint is effective also considering other RFF approaches, in this section, we test our anonymization solution against a different approach considered in the literature, i.e., the use of (raw) IQ samples to train and test a CNN, as considered by [15][13], and [14]. In order to provide a fair comparison among RFF approaches, we consider the ResNet-18 network and re-adapt it to accept as input chunks of $10^{4}$ IQ samples, so to generate a classification for our pool of 10 (ten) transmitters. The size of the chunk has been evaluated as a trade-off between computation overhead and performance.
@@
We consider the dataset collected in the wireless scenario, training on noise-free samples and then testing on increasing values of $\sigma$ while considering only the Gaussian noise. Figure 13 shows the results of our analysis. We observe that even with the smallest noise level ($\sigma=0.01$), the accuracy of RFF via raw IQ samples drops to the random guess (0.1). Figure 13 should be compared with Fig. 11, i.e., same scenario but different methodology (with and without IQ samples pre-processing into images). We observe that both techniques have similar classification performance with noise-free samples ($\sigma=0$), but image pre-processing is confirmed to be more reliable in the presence of noise: when considering $\sigma=0.01$, the accuracy in Fig. 11 (image-based RFF) is higher than the one in Fig. 13 (raw IQ), although not acceptable for the aim of RFF due to the high value of misprediction ($>0.4$). For higher values of $\sigma$, performance across RFF techniques are the same and confirm the effectiveness of HidePrint.
@@
Figure 13: RFF Performance when considering raw IQ samples and different levels ($\sigma$) of Gaussian noise. Standard techniques using raw IQ samples for training and testing are more sensitive to noise injection, thus allowing perfect anonymization even for $\sigma=0.01$.
9 Selective Radio Fingerprint Disclosure
@@
The results shown in Sect. 7 show that the adversary (either $\mathcal{A}_{1}$ or $\mathcal{A}_{2}$) does not have any advantage in adding additional noisy samples to the training process to increase the accuracy during the testing phase. In such a case, as shown in Fig. 5, the adversary would experience the disappearance of the fingerprint when the transmitters add an increasing level of noise. Nevertheless, in this section, we investigate the possibility of performing Selective Radio Fingerprint Disclosure, i.e., allowing a legitimate receiver to be able to execute RFF even when transmitters are adding noise to the signal, thus authenticating the transmitter while avoiding a non-legitimate party (the adversary) from performing the same.
@@
Let us consider a pool of $N$ devices, each of them being able to select a noise level in a pool of $M$ levels. Our analysis shows that, under standard conditions, an adversary cannot infer the identity of the transmitter by guessing it from all the available possibilities $[N\times M]$. In the following, we show that we can achieve such a property by sharing with a legitimate receiver only the current adopted noise level while keeping it secret from the adversary. In this way, a legitimate receiver can authenticate the transmitter, but the adversary cannot.
We consider that the $N$ transmitters in the network are loosely time-synchronized with the receiver, and that the time $t$ is divided into slots of duration $T$, i.e., $t:[t_{0},\ldots,t_{\infty}]$. During each time slot, a transmitter can perform at most one communication [38], choosing the standard deviation $\sigma$ of the noise level ($n_{i}$) to be added to the signal according to Eq. 7.
$n_{i}=H(s_{k}\ |\ t_{i})\mod M,$ (7)
@@
where $H(\circ)$ is a cryptographically secure hash function, $s_{k}$ is the seed (secret to the adversary) preloaded on each transmitter with ID $k$, and finally, $n_{i}$ is the noise level to be used at time slot $t_{i}$. Each transmitter, at each slot $t_{i}$, computes the noise level to be applied to the radio signal according to a pseudo-random sequence that is unknown to the adversary—under the assumption that the transmitter has not been compromised by the adversary, the seed $s_{k}$ is secret to the adversary. The legitimate receiver, aware of the seed $s_{k}$, can obtain the noise level adopted by the transmitter in the current time instant. Under this assumption, the legitimate receiver has a knowledge advantage with respect to the adversary, and can use the appropriate RFF model with respect to the one of the adversary—which is the best that can be adopted under the considered assumptions, i.e., $\mathcal{A}_{1}$ considering all samples. Specifically, in the slot $t_{i}$, the receiver can use a model trained on only 10 classes, i.e., one per transmitter, each for a specific noise pattern. Any other (unauthorized) receivers, unaware of the secret $s_{k}$, is unaware of the noise level used by each transmitter, and has to use a model trained only on noise-free samples, which is more prone to misclassification than the one used by the legitimate receiver (see Fig. 11). Although we acknowledge that such a strategy requires the legitimate receiver to pre-train (${N\choose M}$ models (for each transmitter, all possible noise levels), such training can be done offline. As an example, considering the configuration of this work, the receiver should first pre-train (${10\choose 6}=210$ models in the offline phase. During the later detection phase, the receiver can pick the correct model according to the time slot $t_{i}$, using Eq. 7, and finally test the samples collected from the radio spectrum.
To prove the effectiveness of such a solution against the adversary, we perform Monte-Carlo simulations (100 iterations) by considering a receiver aware of the noise levels considered by the pool of the transmitters. Figure 14 shows the results of our analysis. Our test set consists of (on average) 30 images per transmitter, each image being the result of $10^{5}$ IQ samples. The average accuracy from Fig. 14(a) exceeds 0.96, while we reported the false positive (FPR) and false negative rates (FNR) in Fig. 14(b). The performance depicted by the configuration in Fig. 14 should be compared with those in Fig. 7, which is based on the logic used by the adversary. The knowledge of the current adopted noise (as available on a legitimate receiver) increases the accuracy from about 0.66 to 0.96, thus giving the legitimate receiver the opportunity to reliably authenticate the transmitter, opposite to any entity in the network (adversary) not being aware of the seeds ($s_{k}$).
@@
Figure 14: Selective Radio Fingerprint Disclosure. Only the receiver aware of the transmitter seeds $s_{k}$ can leverage this information to successfully perform RFF.
Let $\delta$ be the knowledge gap between the legitimate receiver and the adversary, and $p$ be the probability that the legitimate receiver correctly classifies the transmitters. We can express the probability $p_{\mathcal{A}}$ that the adversary classifies correctly the transmitters as $p_{\mathcal{A}}=p-\delta$. Given a sequence $w$ of consecutive observations (images generated from a number of IQ samples), any receiver (including unauthorized ones) can make an educated guess to maximize detection accuracy. Indeed, both the receiver and the adversary can collect $w$ observations and then apply a majority voting decision on independent observations to maximize their chances of guessing the actual labels to be assigned to the images (transmitting sources)—under the assumption that the seed $s_{k}$ is not changed during $w$. The probability of getting $v$ successes in $w$ independent trials is given by the binomial distribution as per Eq. 8.
$P(X=v)={w\choose v}p^{v}(1-p)^{w-v}.$ (8)
@@
$P_{succ}=\sum_{v=\lceil\frac{w}{2}\rceil}^{w}{w\choose v}p^{v}(1-p)^{w-v}.$ (9)
@@
We tested our approach via Monte-Carlo simulations considering $p=0.96$ as the probability of successful RFF by the legitimate receiver, in line with the results in Fig. 11. Moreover, we consider an adversary with different knowledge gaps in the range $\delta=\{0.1,0.2,0.3,0.4,0.5\}$. The results of our analysis are shown in Fig. 15, where we highlight with circles markers the probability $P_{succ}$ of successful RFF by the legitimate receiver, considering majority voting over a growing number of rounds (x-axis). We use different symbols for various values of knowledge gap $\delta$. While increasing the number of rounds $w$ makes the outcome of the majority voting more decisive, we observe how $\delta$ gives a significant advantage to the legitimate receiver. Indeed, while $w=6$ allows the legitimate receiver to experience an overwhelming probability of correctly guessing the transmitting sources (greater than 0.99), an adversary with a knowledge gap $\delta=0.3$, i.e., $p_{\mathcal{A}}=0.66$, experiences the same probability of success while adopting the majority voting, i.e., $P_{succ}\approx p_{\mathcal{A}}$.
Finally, we observe that the previous findings can be leveraged by the transmitter to maximize the probability of successful RFF $P_{succ}$ of the legitimate receiver as opposed to the one of non-authorized receivers, e.g., implementing an update of the seed $s_{k}$ every $w=3$ rounds.
@@
Figure 15: Probability of correct authentication of the transmitter ($P_{succ}$) as a function of the number of rounds (time-slots). A legitimate receiver has an advantage (the knowledge of the seed $s_{k}$) that can be exploited to maximize its probability of authenticating the transmitter, while an adversary not aware of $s_{k}$ has a knowledge gap $\delta$ which reduces its chances of performing the same task (device authentication).
10 Discussion and limitations
@@
This work confirms that RFF can be used to perform de-anonymization attacks at the physical layer against radio devices while proposing an effective and efficient solution to prevent it. The core idea of our proposed solution, HidePrint, relies on injecting noise during the modulation process. The injected noise level should be high enough to hide the fingerprint but enough not to affect the quality of the communication link. Experimental results leveraging both wired and wireless links show that the impact of the noise on the quality of the received signal is less than 0.1 dB while achieving a complete disruption of the fingerprint.
We first collected data from a wired link to estimate the real impact of the noise on the fingerprint, excluding on purpose additional distortion and noise phenomena typical of the wireless channel, such as multipath fading. The ability to identify and track a device is hindered by the injection of noise. This can be achieved by considering any noise pattern and a standard deviation of $\sigma\geq 0.02$. We confirmed our findings for wired and wireless links while training on noise-free samples and testing on any other noise level.
We considered two adversary models as a function of the adversary’s capabilities to collect information from the radio devices. We considered different training configurations spanning between noise-free and noisy samples while taking into account either a target device or a pool of devices. All adversary models considered confirm that the injection of noise prevents the identification and tracking of the device.
Finally, the proposed anonymization technique enables a new form of RFF: selective device fingerprinting. A transmitter can control the ability of the receiver to authenticate it (at the physical layer) by resorting to a shared secret. This way, we achieve the implementation of RFF for a controlled subset of receiving entities, i.e., for many but not for all the receivers eavesdropping on the radio spectrum.
@@
Limitations. This work is affected by typical limitations of experimental setups. The number of transmitting devices is limited to 10, and all identification attacks have been made against that reference set. We highlight that a pool of 10 devices is aligned with other reference RFF works [8, 15]. We only considered the BPSK modulation scheme, which is consistently used in many wireless technologies, making our results immediately applicable. Although higher-order modulation schemes, e.g., QPSK and 64-QAM, could be more sensitive to noise, we expect such modulation schemes to be well robust to a degradation of the SNR of only 0.1 dB, shown in Fig. 11 as necessary to achieve successful anonymization. We also limited the type of experiments in the wireless scenario by taking into account a static wireless scenario where the transmitter and the receiver are separated by a few meters, while we simulated propagation over larger distances by considering an attenuator of 30dB. The considered scenarios are the ones where RFF exhibits more reliable performance, so we reasonably expect our solution to be robust also in more challenging conditions. Finally, we considered only ResNet-18 and specific configuration parameters. Although ResNet-18 is well known to guarantee the best trade-off in terms of accuracy and training time, we plan to test our solution with other CNN while considering different parameters and pre-processing configurations.
11 Conclusion and Future Work
@@
We have presented HidePrint, a technique to prevent RFF by injecting random noise in the emitted RF signal without affecting communication quality, thus mitigating physical layer attacks aimed at identifying and tracking radio devices. We have tested our solution against state-of-the-art classification techniques, different adversary models, and scenarios. We have proven that the injection of a minimal amount of noise in the transmitted signal does not affect the quality of the link but completely hides the fingerprint from the receiver. Moreover, we have discussed the use of HidePrint for selective radio fingerprint disclosure, enabling only a subset of legitimate receivers to perform RFF while excluding other (unauthorized) ones. Future works include testing our solution against different classifiers while considering different deployment scenarios.

### v2

@@ 8. Comparison
The RFF methodology used in the previous section to validate HidePrint involves the pre-processing of IQ samples into images, as described in Sect. 6. As shown by the authors in (alhazbi2023_acsac), this strategy is effective in mitigating the impact of noise on the RFF process. To provide further insights and show that HidePrint is effective also considering other RFF approaches, in this section, we test our anonymization solution against a different approach considered in the literature, i.e., the use of (raw) IQ samples to train and test a CNN, as considered by (liu2023_wcl)(sun2022_rs), and (lu2024_arxiv). We highlight that for image-based pre-processing, we resort to the implementation of CNN in MatLab that cannot be used out of the box, but it requires some modification to the input and output layers. It is worth mentioning the choice of using an already designed (and trained) network, despite designing and training it from scratch. Indeed, considering a self-designed network makes the comparison with the state-of-the-art even more challenging, since all the contributions adopting such a strategy are not consistent in the design of the neural network. Therefore, to provide a fair comparison, we consider ResNet-18 and re-adapt it to accept as input chunks of $10^{4}$ IQ samples. The size of the chunk has been evaluated as a trade-off between computation overhead and performance.
@@
We consider the dataset collected in the wireless scenario, training on noise-free samples and then testing on increasing values of $\sigma$, with Gaussian noise. Figure 13 shows our results. We observe that even with the smallest noise level ($\sigma=0.01$), the accuracy of RFF via raw IQ samples drops to the random guess (0.1). Figure 13 should be compared with Fig. 11, i.e., same scenario but different methodology (with and without pre-processing into images). We observe that both techniques have similar performance with noise-free samples ($\sigma=0$), but image pre-processing is confirmed to be more reliable with noise: when considering $\sigma=0.01$, the accuracy in Fig. 11 (image-based RFF) is higher than in Fig. 13 (raw IQ), although not acceptable for RFF due to too high mispredictions (>0.4>0.4). For higher values of $\sigma$, performance across RFF techniques are the same and confirm the effectiveness of our solution.
@@
Figure 13. RFF Performance when considering raw IQ samples and different levels ($\sigma$) of Gaussian noise. Standard techniques using raw IQ samples for training and testing are more sensitive to noise injection, thus allowing perfect anonymization even for $\sigma=0.01$.
9. Selective Fingerprint Disclosure
@@
Our results (Sect. 7) show that the adversary (both $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$) does not gain any advantage in adding additional noisy samples to the training process. In such scenarios, as shown in Fig. 5, the adversary experiences the disappearance of the fingerprint when the transmitters add an increasing level of noise. Nevertheless, in this section, we investigate the possibility of performing Selective Radio Fingerprint Disclosure, i.e., allowing a legitimate receiver to successfully perform RFF even when transmitters are adding noise to the signal, while avoiding the same for illegitimate parties (adversary).
@@
We consider a pool of $N$ devices, each of them being able to select a noise level in a pool of $M$ levels. Our following analysis shows that, under standard conditions, an adversary cannot infer the identity of the transmitter from all the available possibilities $N\times M]$. In the following, we show that we can achieve such a property by sharing with a legitimate receiver only the noise level currently adopted while keeping it secret from the adversary. In this way, a legitimate receiver can authenticate the transmitter, but the adversary cannot. We consider that the $N$ transmitters are loosely time-synchronized with the receiver, and that the time $t$ is divided into slots of duration $T$, i.e., $t:[t_{0},\ldots,t_{\infty}]$. During each time slot, a transmitter can perform at most one communication (dipietro_2017_cn), choosing the standard deviation $\sigma$ of the noise level ($n_{i}$) to be added to the signal via Eq. 7.
(7) $n_{i}=H(s_{k}\ |\ t_{i})\mod M,$
@@
being $H(\circ)$ a cryptographically secure hash function, $s_{k}$ the seed (secret to the adversary) preloaded on each transmitter with ID $k$, and $n_{i}$ the noise level to be used at time slot $t_{i}$. Each transmitter, at each slot $t_{i}$, computes the noise level to be applied to the radio signal according to a pseudo-random sequence that is unknown to the adversary— the seed $s_{k}$ is secret to the adversary. The legitimate receiver, aware of the seed $s_{k}$, can compute the noise level adopted by the transmitter in the current time slot. Under this assumption, the legitimate receiver has a knowledge advantage with respect to the adversary, and can use the appropriate RFF model with respect to the one of the adversary—which is the best that can be adopted under the considered assumptions, i.e., $\mathcal{A}_{1}$ considering all samples. Specifically, in the slot $t_{i}$, the receiver can use a model trained on only 10 classes, i.e., one per transmitter, each for a specific noise pattern. Any other (unauthorized) receivers, unaware of the secret $s_{k}$, do not know the noise level used by each transmitter, and has to use a model trained only on noise-free samples, which is more prone to misclassification than the one used by the legitimate receiver (see Fig. 11). We acknowledge that such a strategy requires the legitimate receiver to pre-train $N\choose M$} models (for each transmitter, all possible noise levels). However, the training can be done offline. As an example, considering the configuration of this work, the receiver should first pre-train ${10\choose 6}=210$ models in the offline phase. At runtime, the receiver can pick the correct model according to the time slot $t_{i}$, using Eq. 7, and finally test the samples collected from the radio spectrum. To prove the effectiveness of such a solution against the adversary, we perform Monte-Carlo simulations (100 iterations) by considering a receiver aware of the noise levels considered by the pool of the transmitters. Figure 14 shows the results of our analysis. Our test set consists of (on average) 30 images per transmitter, each image being the result of $10^{5}$ IQ samples. The average accuracy from Fig. 14(a) exceeds 0.96, while we reported the FPR and FNR in Fig. 14(b). The performance in Fig. 14 should be compared with those in Fig. 7, which is based on the logic used by the adversary. The knowledge of the noise currently adopted (available on a legitimate receiver) increases the accuracy from 0.66 to 0.96, allowing the receiver to reliably authenticate the transmitter.
@@
Figure 14. Selective Radio Fingerprint Disclosure. Only the receiver aware of the seeds $s_{k}$ can successfully perform RFF.
Let $\delta$ be the knowledge gap between the legitimate receiver and the adversary, and $p$ be the probability that the legitimate receiver correctly classifies the transmitters. We can express the probability $p_{\mathcal{A}}$ that the adversary classifies correctly the transmitters as $p_{\mathcal{A}}=p-\delta$. Given a sequence $w$ of consecutive observations (images generated from a number of IQ samples), any receiver (including unauthorized ones) can make an educated guess to maximize detection accuracy. Indeed, both the receiver and the adversary can collect $w$ observations and then apply a majority voting decision on independent observations to maximize their chances of guessing the actual labels to be assigned to the images (transmitting sources)—under the assumption that the seed $s_{k}$ is not changed during $w$. The probability of getting $v$ successes in $w$ independent trials is given by the binomial distribution according to Eq. 8.
(8) $P(X=v)={w\choose v}p^{v}(1-p)^{w-v}.$
@@
(9) $P_{succ}=\sum_{v=\lceil\frac{w}{2}\rceil}^{w}{w\choose v}p^{v}(1-p)^{w-v}.$
@@
We tested our approach via Monte-Carlo simulations considering $p=0.96$ as the probability of successful RFF by the legitimate receiver, as previously obtained and reported in Fig. 11. Moreover, we consider an adversary with different knowledge gaps in the range $\delta=\{0.1,0.2,0.3,0.4,0.5\}$. Our results are shown in Fig. 15, where we highlight with circle markers the probability $P_{succ}$ of successful RFF by the legitimate receiver, considering majority voting over a growing number of rounds (x-axis). We use different symbols for various values of the knowledge gap $\delta$. While increasing the number of rounds $w$ makes the outcome of the majority voting more decisive, we observe how $\delta$ gives a significant advantage to the legitimate receiver. The value $w=6$ allows the legitimate receiver to experience an overwhelming probability of correctly guessing the transmitter (greater than 0.99). However, an adversary with a knowledge gap $\delta=0.3$, i.e., $p_{\mathcal{A}}=0.66$, experiences the same probability of success while adopting the majority voting, i.e., $P_{succ}\approx p_{\mathcal{A}}$. Finally, note that the transmitter can maximize the probability of successful RFF $P_{succ}$ of the legitimate receiver as opposed to the one of non-authorized receivers, e.g., updating the seed $s_{k}$ every $w=3$ rounds.
@@
Figure 15. Probability of correct transmitter authentication ($P_{succ}$) with various rounds (time-slots). A legitimate receiver has an advantage (the knowledge of the seed $s_{k}$) that can be exploited to maximize its probability of authenticating the transmitter, while an adversary not aware of $s_{k}$ has a knowledge gap $\delta$ which reduces its chances of performing the same task (device authentication).
10. Discussion
@@
Main Findings. Our work confirms that RFF can be used to perform de-anonymization attacks at the physical layer against radio devices while proposing an effective and efficient solution to prevent it. The core idea of our proposed solution, HidePrint, relies on injecting noise during the modulation process. The injected noise level should be, at the same time, high enough to hide the fingerprint and low enough not to affect the quality of the communication link. Experimental results leveraging both wired and wireless links show that the impact of the noise on the quality of the received signal is less than 0.1 dB while achieving a complete disruption of the fingerprint. We first collected data from a wired link to estimate the real impact of the noise on the fingerprint, excluding, on purpose, additional distortion and noise phenomena typical of the wireless channel, such as multipath fading. The ability to identify and track a device is hindered by the injection of noise. This can be achieved by considering any noise pattern and a standard deviation of $\sigma\geq 0.02$. We confirmed our findings for wired and wireless links while training on noise-free samples and testing on any other noise level. We also confirmed the robustness of HidePrint considering two adversary models as a function of the adversary’s capabilities to collect information from the radio devices. We considered different training configurations spanning between noise-free and noisy samples while taking into account either a target device or a pool of devices. All the considered adversary models confirm that the injection of noise prevents the identification and tracking of the device. Finally, we proposed a new form of RFF: selective device fingerprinting. A transmitter can control the ability of the receiver to authenticate it (at the physical layer) by resorting to a shared secret. We can thus achieve effective RFF for a controlled subset of receiving entities, i.e., for many but not for all receivers eavesdropping on the radio spectrum.
Robustness to Channel Equalization Techniques. An orthogonal approach that the adversary could use to improve the accuracy of transmitter identification is to use channel equalization techniques. These techniques have shown a remarkable capability to correct transmission errors when the noise is introduced by the wireless channel. Note that channel equalization techniques work by assuming that the statistical properties of the wireless channel are predictable and stable over short periods, and that they do not change too quickly (haykin2008_book). In our context, the transmitter using HidePrint continuously switches the noise type and the standard deviation of the injected noise ($\sigma$). This behavior breaks, by definition, any assumptions about the predictability and stability of the noise added to the signal, making those techniques unreliable. Moreover, channel equalization techniques have been conceived to recover the symbols (and thus, the bits) transmitted by the receiver, in the presence of distortions (noise and multipath). They do not recover the original value of the IQ sample before noise addition. Thus, the application of those techniques reshapes (by design) the radio fingerprint experienced by the receiver.
@@
Limitations. We consider a limited set of 10 transmitting devices, in accordance with other reference RFF works (alhazbi2023_acsac; lu2024_arxiv). We only considered the BPSK modulation scheme, which is consistently used in many wireless technologies, making our results immediately applicable. We did not apply our solution to other higher-order modulation schemes. Although schemes like QPSK and 64-QAM could be more sensitive to noise, such modulation schemes are well robust to a degradation of the SNR of only 0.1 dB (see Fig. 11). We also limited the type of experiments in the wireless scenario by considering a static wireless scenario, while we modelled propagation over larger distances by considering various SNR levels and an attenuator of 30dB. These scenarios are the ones where RFF exhibits more reliable performance, so this methodology does not affect our findings. Finally, we considered only the image-based RFF solutions based on ResNet-18 and AEs, based on relevant recent literature (oligeri2023tifs)(alhazbi2023_acsac), and RFF approaches based on raw IQ samples (liu2023_wcl)(sun2022_rs)(lu2024_arxiv). In our future work, we will extend our analysis with further classifiers.
11. Conclusion and Future Work
@@
We have presented HidePrint, a technique to prevent hostile RFF by injecting random noise in the emitted RF signal without affecting communication quality, thus mitigating physical layer attacks aimed at identifying and tracking radio devices. We have tested our solution against state-of-the-art classification techniques, different adversary models, and scenarios. We have proven that the injection of a minimal amount of noise in the transmitted signal does not affect the quality of the link but completely hides the fingerprint from the receiver. Moreover, we have discussed the use of HidePrint for selective radio fingerprint disclosure, enabling only a subset of legitimate receivers to perform RFF while excluding unauthorized ones. Future works include testing our solution against different classifiers while considering different scenarios.